OpenStudy (anonymous):

: How long does it take to send a 1.5 Kbyte file over a 2.9 Mbytes/s channel that is 100.0 km long? Assume speed of light in that channel is 2.0x10 8 m/s. Give your answer in ms. , , @Physics

OpenStudy (anonymous):

2.0 x 10^8 m/s *correction

OpenStudy (stormfire1):

hmm...I have to think about this for a min

OpenStudy (anonymous):

sure np.

OpenStudy (stormfire1):

Now I'm questioning how I did the previous one. One min..

OpenStudy (anonymous):

previous one was right, at least according to the homework console. also, it says that the answer to this one is 1 ms.

OpenStudy (stormfire1):

I thought you said the first answer was 5 x 10^-1?

OpenStudy (anonymous):

oh sorry I take that previous comment back lol. it really did say 5x10^-1 ms.

OpenStudy (stormfire1):

After some more thought, I believe I was wrong about the first one. The speed of light doesn't matter in either of these. The limiting factor is the channel which is much slower

OpenStudy (anonymous):

got 1.017 ms here, but homework console takes significant digits into account, so 1 ms is fine. please confirm.

OpenStudy (stormfire1):

ok...just to be sure, the channel is in Mbps or Mbytes/sec?

OpenStudy (stormfire1):

bits or bytes?

OpenStudy (anonymous):

Mbytes/sec :D

OpenStudy (anonymous):

all are bytes here

OpenStudy (stormfire1):

Yes, I got .0010172414 s or 1 ms

OpenStudy (anonymous):

thanks! :D

OpenStudy (stormfire1):

np :)

OpenStudy (anonymous):

\[1.5kB \times \frac{1MB}{1024kB} \times \frac{1s}{2.9MB} \times \frac{1000ms}{1s}\]That's how long it takes to send the data packet.\[100km \times \frac{1000m}{1km} \times \frac{1s}{2\times10^{8}m} \times \frac{1000ms}{1s}\]That's the propagation time for the channel. I think they're probably going for the sum of these two times.

OpenStudy (anonymous):

yep. that's right. thanks too! :D

OpenStudy (stormfire1):

I'm no CompSci major so I just used the formula :)

OpenStudy (anonymous):

Dimensional analysis. Learned it in high school chemistry. It never fails.

OpenStudy (stormfire1):

I normally do it on paper...it's just a pain to use the equation editor here so I skip it a lot

OpenStudy (anonymous):

the key was adding the two times.. i tried dimensional analysis then wasn't sure what to do after that haha

OpenStudy (stormfire1):

Sadly, I had to wait until I was 40 to learn it...it was never taught in any of my High School classes :)

OpenStudy (anonymous):

Yeah, I got stuck when trying to put that conversion info into the first expression, and that made me realize it really wasn't part of the first quantity, and that there were actually two quantities.